OpenStudy (anonymous):

y = 2x + 9 y = 4x + 4 a. (2.5, 14) b. (4, –2.5) c. (–2.5, 4) d. (–2.5, –6)

OpenStudy (anonymous):

okay i got you jackson

OpenStudy (anonymous):

When you see an equals sign, it means that the thing on the left is the EXACT SAME THING as the thing as the right. Whenever you see the thing on the left, you can replace it with the thing on the right... because they're the same thing. Substitution uses this to make an easier problem. In this example, you must solve for two variables, x and y. Using substitution, you can solve for just one, and then substituting again, find the other. y = 4x - 7 y = 2x + 9 Replacing first y with 2x + 9 2x + 9 = 4x - 7 Subtract 9. 2x = 4x - 16 Subtract 4x. -2x = -16 Divide by -2/ x = 8 Substituting the new value of x in the first equation, y = 4x - 7 y = 4*8 - 7 y = 32 - 7 = 25 x = 8, y = 25 (8, 25) Plug the answer back in to make sure everything works.

OpenStudy (anonymous):

y = 2x + 9 y = 4x + 4 2x + 9 = 4x + 4 -- subtract 4x 2x - 4x + 9 = 4 -- subtract 9 2x - 4x = 4 - 9 -- combine like terms -2x = - 5 x = 5/2 x = 2.5 y = 2x + 9 y = 2(2.5) + 9 y = 5 + 9 y = 14 x = 2.5 and y = 14

OpenStudy (anonymous):

thank you can you help me pleasee

OpenStudy (anonymous):

Two groups of students order burritos and tacos at a local restaurant. One order of 3 burritos and 4 tacos costs $11.33. The other order of 9 burritos and 5 tacos costs $23.56. What is the cost of a taco? a. $1.49 c. $0.99 b. $1.79 d. $0.89

OpenStudy (anonymous):

lets say that : the cost of one burrito= x the cost of one taco= y then, 3x+4y=11.33 9x+5y=23.56 that was the system the second question: if u multiply the first equation with -3 it will be -9x-12y=-33.99 9x+5y=23.56 -7y=-10.43 y=1.49 by plugging y in the first equation u will have x=1.79 so the cost of one burrito= 1.79 the cost of one taco= 1.49